OpenStudy (anonymous):

A jar contains 20 yellow marbles, 55 green marbles, and 25 purple marbles. You pick one marble from the jar at random. What is the theoretical probability of picking a green marbel? A. 55% B. 45% C. 25% D. 20%

OpenStudy (anonymous):

total probability 55+25+20=100 p(green)=55/100 * 100 =55%